OpenStudy (anonymous):

For a triangle with sides a, b and c, what is the range of the length of side c? a = 418, b = 222, c =

OpenStudy (anonymous):

a^2+c^2=b^2 now you switch it around c^2-a^2=b^2 i think ur doing pythagorium